# Variational Adaptive Gaussian Decomposition: Scalable Quadrature-Free Time-Sliced Thawed Gaussian Dynamics

### 11 Mar 2026 v1

@@ Abstract
Time-slicing has emerged as a strategy for incorporating semiclassical propagation into real-time path integral formulation and recovering full quantum mechanical dynamics. A central step is the decomposition of a time-evolved wave function into a superposition of Gaussian wave packets. Here we introduce a quadrature-free variational framework for Gaussian wave packet decomposition, reformulating it as an optimization problem in which the parameters of Gaussian wave packets are chosen to maximize the overlap with the time-evolving wave function. An autoencoder-decoder neural network is used for this optimization, with the representation being adaptively reoptimized during propagation. Each wave packet in this decomposition represents a localized patch of the underlying semiclassical manifold, while retaining full correlations between all degrees of freedom. This variational adaptive Gaussian decomposition (VAGD) approach yields a compact Gaussian expansion, providing a scalable route to time-sliced semiclassical quantum dynamics. While general, applying VAGD to facilitate time-slicing of thawed Gaussian approximation (TGA) simulation allows a route to improving the semiclassical result to the full quantum mechanical result in a systematic manner.
@@ I Introduction
Much of chemistry can be formulated in terms of the time-evolution of systems governed by quantum mechanics. However, simulating the time-dependent Schrödinger equation poses significant computational challenges as the complexity grows exponentially with the number of degrees of freedom. Formally exact methods for wave-function propagation such as the family of multiconfigurational time-dependent Hartree methods ((ML-)MCTDH) [21, 20] are able to handle the systems with up to a few tens of degrees of freedom. However, despite their remarkable efficiency, these formally exact methods are limited by the growing Hilbert space dimensionality. Consequently, approximate approaches that exploit classical insights prove to be very powerful. Owing to Bohr’s Correspondence Principle and the relatively large masses of atoms and molecules, these semiclassical schemes of propagation have, over the years, emerged as a framework of accurate approximations to the full quantum dynamics [22]. These methods also have the added advantage that they can be applied to arbitrary potential energy landscapes in an approximate manner.
@@
Among the many semiclassical techniques that have been developed, two prominent classes are those based on stationary-phase approximation of the path integral [25, 16, 23] and the thawed Gaussian approximation [14, 15, 12, 13, 26] (TGA). TGA is a particularly attractive approximate technique that describes the motion of a Gaussian wave packet (GWP) whose width and orientation can evolve dynamically on an arbitrary potential energy surface. Simpler versions of this method have even been applied to study systems in an ab initio manner [3, 2, 1, 4, 5]. While extremely attractive, the accuracy of these semiclassical methods typically deteriorates at long propagation times. A crucial approach to improving this is the idea of re-expanding the time-dependent wave-function into a superposition of Gaussian wave packets (GWPs) or time-slicing [11, 6, 17]. This Gaussian decomposition is done at intervals where the semiclassical propagation falters. Similar ideas are at the heart of Gaussian MCTDH [8], its system-bath versions [9, 7, 24], and the related multi-level methods [19]
@@
In improving semiclassical propagation at long times, there are two challenges, therefore, that must be addressed: (1) the efficiency of the algorithm for time-slicing for a multidimensional system; (2) the number of resultant Gaussians in the decomposed wave function. The first is a challenge because the procedure for decomposing a wave function into many Gaussians involves multidimensional integrations. Although Monte Carlo-based methods are often the most practical methods for handling multidimensional integrals, they become inefficient when the integrand has both positive and negative parts. In such cases, phase cancellation leads to an exponential growth of variance, giving rise to the Monte Carlo “sign problem”, which becomes increasingly severe with dimensionality. Moreover, even wave functions that are initially real acquire complex phases during time-evolution, which leads to the infamous “dynamical sign-problem”.
@@
Kong et al. [17] showed that using a Husimi transform expansion one can solve even the strongly correlated problem of dynamics in a two-dimensional double well potential at full quantum levels of accuracy using the time-sliced TGA (TSTG). However, despite its efficiency, this work uses a quadrature-based implementation of fast Gaussian transforms. While this TSTG avoided the Monte Carlo sign problem through an innovative use of fast Gaussian transforms, it is still inherently quadrature-based. Consequently, the number of Gaussians, and therefore classical trajectories, grows exponentially with the system dimensionality. This scaling severely limits the practical applicability of TSTG in combination with ab initio molecular dynamics.
@@
In this work, we propose an alternative approach to Gaussian wave function decomposition that circumvents both the problems. The key insight is to reformulate the time-slicing as an optimization problem in which the parameters of the GWPs constituting the decomposed output are determined variationally, using an autoencoder-decoder neural network, to maximize the overlap with the input wave function. This variational adaptive Gaussian decomposition (VAGD) method is quadrature-free, which allows for multidimensional applications, and produces an optimally compact representation, minimizing the number of classical trajectories generated. VAGD is general, and can be used in conjunction with any semiclassical propagation to enable time-slicing.
@@ II.1 Time-Sliced Thawed Gaussian
where using the local harmonic approximation, the “position” and “momenta” coordinates, $\vec{q}_{t}$ and $\vec{p}_{t}$ respectively, satisfy the classical equations of motion,
@@
While exact for harmonic potentials, owing to the impact of the higher-order derivatives of the anharmonic potentials, the quality of TGA deteriorates with time even when the initial state is a Gaussian. As outlined originally in Ref. 11, this deterioration can be alleviated by essentially using the semiclassical propagator as an ingredient for a path integral simulation. The essential idea is to propagate with TGA till the prescription starts to break down, at which point, the wave function is re-expressed as a superposition of localized GWPs. Each of these wavepackets are then individually propagated once again till the next decomposition step. The frequency of the decomposition step is a convergence parameter, and denoted by $N_{\text{seg}}$.
@@ II.2 Variational Adaptive Gaussian Decomposition (VAGD)
We introduce a variational approach for constructing an optimized Gaussian representation of an arbitrary input wave function. In the context of VAGD with TGA, the input wave function is written as a sum of $N_{\text{inp}}$ Gaussians,
[…]
@@
With this perspective, it is not necessary to train a generically applicable neural network. Instead, the network is optimized specifically for each wave function under consideration. In applications where the propagation of classical trajectories is the computational bottleneck, the additional cost associated with the optimization is negligible. In fact, it is even beneficial to increase the time for a proper optimization in an attempt to decrease the number of classical trajectories. Moreover, because the network is used as a variational optimizer rather than a predictive model, the usual challenges associated with neural networks, such as the construction of representative training sets and concerns about overfitting, are not relevant here. In this sense, the variational optimization adaptively identifies the set of Gaussian wave packets that best cover the relevant region of phase space explored by the input wave function.
Finally, we discuss how $N_{\text{out}}$ is determined. The optimal representation is defined by the minimum number of Gaussians used, or minimum $N_{\text{out}}$, required to achieve $F_{\text{thresh}}$, essential for reducing the number of classical trajectories. However, as end-users, we only set $K$, which is an upper limit to $N_{\text{out}}$. As an initial guess, the algorithm sets $N_{\text{out}}=N_{\text{inp}}$. If upon optimization, the threshold fidelity, $F_{\text{thresh}}$, is not achieved, then the value of $N_{\text{out}}$ is increased by a user-specified step size $\Delta N$ and the network is reoptimized. The process is repeated till either the threshold fidelity is successfully achieved or the upper limit of $K$ is reached. We refer to this procedure as an adaptive basis expansion technique, which ensures an optimal Gaussian decomposition.
@@ II.2.1 Encoding of Parameters
Now the interpretation of the output is largely similar, with the $(\vec{q},\vec{p})$ and $\gamma$, retained. The main difference comes in the $A$ matrix: the real part of $A$, $A^{\text{R}}$, must be symmetric, but the imaginary part of $A$, $A^{\text{I}}$, which encodes the wave packet width, must be positive definite. Directly optimizing $A^{\text{I}}$ during training does not guarantee this property. To ensure positive definiteness by construction, we optimize a lower triangular matrix, $L$, with positive diagonal elements for each Gaussian wave packet. The width matrix of the $k$th Gaussian is finally reconstructed using a Cholesky decomposition,
$A_{k}=A_{k}^{\text{R}}+iL_{k}L_{k}^{T}.$ (10)
@@ II.2.2 Numerical Stabilization Strategies
VAGD, as defined so far, allows for arbitrarily broad Gaussians to be incorporated in the decomposition if required. While broad Gaussians help minimize the number of trajectories, the local harmonic approximation is violated sooner for them than for the narrow wave packets, necessitating a more frequent decomposition to maintain accuracy. This issue can be mitigated by putting a dimension-specific, user-defined cap on the width of each Gaussian, which effectively sets a lower limit on the diagonal values of $L_{k}$ defined in Eq. (10). We refer to the minimum inverse width parameter as $W_{\text{min}}$, which is the lower bound on diagonal elements of A matrix, effectively putting an upper bound on spatial widths of the Gaussians.
A direct implementation of a VAGD decomposition step as described till now through the duration of propagation starts from a randomly initialized network, which increases the number of optimization cycles, or epochs, required for convergence. Additionally, the position and the width of the wave function can fluctuate significantly during time-evolution, which makes the numerical optimization of the neural network difficult. We will see that these two difficulties are closely related.
@@
This regularization consists of two preprocessing steps to ensure a faster training with the warm-start ansatz. First, the wave function is recentered to remove its average position. Second, the coordinate axes are rescaled so that the wave function has variances comparable to the initial state. The neural network is trained on this preprocessed wave function. After optimization, the regularization transformation is undone to recover the physical representation. This combination of regularization and warm-start initialization will be referred to as the regularized warm-start training.
@@ III.1 Morse Potential
We start our exploration with the dynamics of wave packets in Morse potentials. In Sec. III.1.1, we explore the convergence of VAGD-TGA to the full quantum results under different levels of anharmonicity.
@@
where $D_{e}$ is the dissociation energy, $a$ is the width of the well, and $x_{\text{eq}}$ is the equilibrium bond distance. The harmonic frequency at the bottom of the well is given by $\omega=a\sqrt{2D_{e}}$. It is often more convenient to think in terms of the harmonic frequency $\omega$ and an anharmonicity factor $\chi=\frac{\omega}{4D_{e}}$[1], which is zero for harmonic potentials. The parameters, $D_{e}$ and $a$ can be simply obtained by inverting the relations.
In Fig. 2, we show the overlap of the wave function evolved using the traditional TGA with the numerically exact SOFT simulations for different values of $\chi$. Following the examples in Ref. 1, the equilibrium bond length $x_{\text{eq}}=$20.95\text{\,}\mathrm{\text{ $a_{0}$ }}$\$ and $\omega=$900\text{\,}{\mathrm{cm}}^{-1}$\$. The initial wave packet is taken to be a real Gaussian centered at $q_{0}=0$, with $A=-\frac{\omega_{0}}{\hbar}$, where $\omega_{0}=$1000\text{\,}{\mathrm{cm}}^{-1}$\$. The simulations were run with a fixed time-step of $\Delta t=$0.053\text{\,}\mathrm{fs}$\$. Notice that, consistent with our expectations, in Fig. 2, the quality of TGA deteriorates with time, and this deterioration is faster for larger values of $\chi$.
@@
Now, we explore our VAGD method. We perform the variational decomposition every 50 steps (segment length $\tau=50\,\Delta t$). At each decomposition boundary, the reconstructed state is required to satisfy $F_{\text{thresh}}=0.9995$ with the propagated. wavefunction. The $W_{min}$ was set to 5.0.
Consider the case of $\chi=0.005$ as a first example. In Fig. 3, we examine the improvement of VAGD with increasing $K$. Notice that there is a clear and systematic improvement of the overlap with the SOFT result with the maximum number of GWPs allowed, $K$. This improvement is most clearly observed at longer times, as the quality of the semiclassical approximation gets worse. By $K\geq 6$, the VAGD-TGA becomes practically identical to SOFT. For this example, we push the limits and claim convergence at $K=8$.
Figure 4: Overlap of VAGD-TGA wave function with SOFT solution (solid, left $y$-axis) and number of VAGD trajectories (dashed, right $y$-axis) of 1D-Morse oscillator for different values of $\chi$
On increasing anharmonicity, while $\chi=0.01$ ran with the same decomposition frequency, $\chi=0.02$ required more frequent decomposition at every 25 steps. We show the overlap and the number of trajectories for the converged simulation in Fig. 4. The convergence gets more difficult with anharmonicity. One way to think about this is greater the anharmonicity, thinner are the GWPs required to maintain the validity of the local harmonic approximation over the same segment length. If the GWPs involved in the resolution of identity are thinner, a higher number of them would be required, thereby necessitating a larger $K$ and a larger value of $N_{\text{traj}}(t)$.
Also notice that the number of trajectories, $N_{\text{traj}}(t)$, grows with time demonstrating the build-up of the effects of anharmonicity and deviation from the basic thawed Gaussian ansatz. However, what is also extremely heartening is that the number of trajectories required is relatively small in all the cases. This becomes important especially when running ab initio dynamics where trading the cost of a better decomposition for fewer number of trajectories can prove to be extremely lucrative. Moreover, depending on the physics of the problem at hand, the actual number of trajectories may be substantially less than the upper limit $K$ provided. This is another benefit of the adaptive algorithm. It is also to be noted that checking the overlap with the numerically exact SOFT results is by far the most stringent test of the correctness of the method. Other observables, like correlation functions, that are crucial for spectra, would converge with far fewer trajectories.

### 12 Mar 2026 v2

@@ Abstract
Time-slicing has emerged as a strategy for incorporating semiclassical propagation into real-time path integral formulation and recovering full quantum dynamics. A central step is the decomposition of a time-evolved wave function into a superposition of Gaussian wave packets (GWPs). Here we introduce a quadrature-free variational framework for GWP decomposition, reformulating it as an optimization problem in which the GWP parameters are chosen to maximize the overlap with the time-evolving wave function. An autoencoder-decoder neural network is used for this optimization, with the representation being adaptively reoptimized during propagation. Each wave packet in this decomposition represents a localized patch of the underlying semiclassical manifold, while retaining full correlations between all degrees of freedom. This variational adaptive Gaussian decomposition (VAGD) approach yields a compact Gaussian expansion, providing a scalable route to time-sliced semiclassical quantum dynamics. While general, applying VAGD to facilitate time-slicing of thawed Gaussian approximation (TGA) allows a route to improving the semiclassical treatment to the full quantum mechanical result in a systematic manner.
@@ I Introduction
Much of chemistry can be formulated in terms of the time-evolution of systems governed by quantum mechanics. However, simulating the time-dependent Schrödinger equation poses significant computational challenges as the complexity grows exponentially with the number of degrees of freedom. Formally exact methods for wave-function propagation such as the family of multiconfigurational time-dependent Hartree methods ((ML-)MCTDH) [21, 20] are able to handle the systems with up to a few tens of degrees of freedom. However, despite their remarkable efficiency, these methods are limited by the growing Hilbert space dimensionality. Consequently, approximate approaches that exploit classical insights prove to be very powerful. Owing to Bohr’s Correspondence Principle and the relatively large masses of atoms and molecules, these semiclassical schemes of propagation have, over the years, emerged as a framework of accurate approximations to the full quantum dynamics [22]. These methods also have the added advantage that they can be applied to arbitrary potential energy landscapes in an approximate manner.
@@
Among the many semiclassical techniques that have been developed, two prominent classes are those based on stationary-phase approximation of the path integral [25, 16, 23] and the thawed Gaussian approximation [14, 15, 12, 13, 26] (TGA). TGA is a particularly attractive approximate technique that describes the motion of a Gaussian wave packet (GWP) whose width and orientation can evolve dynamically on an arbitrary potential energy surface. Simpler versions of this method have even been applied to study molecular systems in an ab initio manner [3, 2, 1, 4, 5]. While extremely attractive, the accuracy of these semiclassical methods typically deteriorates at long propagation times. A crucial approach to improving this is the idea of re-expanding the time-dependent wave-function into a superposition of Gaussian wave packets (GWPs) or time-slicing [11, 6, 17]. This Gaussian decomposition is done at intervals beyond which the semiclassical propagation falters. Similar ideas are at the heart of Gaussian MCTDH [8], its system-bath versions [9, 7, 24], and the related multi-level methods [19].
@@
In improving semiclassical propagation at long times, there are two challenges, therefore, that must be addressed: (1) the efficiency of the algorithm for time-slicing for a multidimensional system; (2) the number of resultant Gaussians in the decomposed wave function. The first is a challenge because the procedure for decomposing a wave function into many Gaussians involves multidimensional integrals. Although Monte Carlo-based methods are often the most practical methods for handling such multidimensional integrals, they become inefficient when the integrand has both positive and negative parts. In such cases, phase cancellation leads to an exponential growth of variance, giving rise to the infamous Monte Carlo “sign problem”, which becomes increasingly severe with dimensionality. Moreover, even wave functions that are initially real acquire complex phases during time-evolution leading to the “dynamical sign-problem.”
@@
Kong et al. [17] showed that using a Husimi transform expansion one can solve even the strongly correlated problem of dynamics in a two-dimensional double well potential at full quantum levels of accuracy using the time-sliced TGA (TSTG). While TSTG avoids the Monte Carlo sign problem through an innovative use of fast Gaussian transforms, it is still inherently quadrature-based. Consequently, the number of Gaussians, and therefore classical trajectories, grows exponentially with the system dimensionality. This scaling severely limits the practical applicability of TSTG in combination with ab initio molecular dynamics.
@@
In this work, we propose an alternative approach to Gaussian wave function decomposition that circumvents both the problems. The key insight is to reformulate time-slicing as an optimization problem in which the parameters of the GWPs constituting the decomposed output are determined variationally, using an autoencoder-decoder neural network, to maximize the overlap with the input wave function. This variational adaptive Gaussian decomposition (VAGD) method is quadrature-free, which allows for multidimensional applications, and produces an optimally compact representation, minimizing the number of classical trajectories generated. VAGD is general, and can be used in conjunction with any semiclassical propagation to enable time-slicing.
@@ II.1 Time-Sliced Thawed Gaussian
where under the local harmonic approximation, the “position” and “momenta” coordinates, $\vec{q}_{t}$ and $\vec{p}_{t}$ respectively, satisfy the classical equations of motion,
@@
While exact for harmonic potentials, owing to the impact of the non-zero higher-order derivatives of the anharmonic potentials, the quality of TGA deteriorates with time even when the initial state is a Gaussian. As outlined originally in Ref. 11, this deterioration can be alleviated by essentially using the semiclassical propagator as an ingredient for a path integral simulation. The essential idea is to propagate with TGA till the prescription starts to break down, at which point, the wave function is re-expressed as a superposition of localized GWPs. Each of these wavepackets are then individually propagated once again till the next decomposition step. The number of time-steps between consecutive decomposition steps is a convergence parameter, and denoted by $N_{\text{seg}}$.
@@ II.2 Variational Adaptive Gaussian Decomposition (VAGD)
With this perspective, it is not necessary to train a generically applicable neural network. Instead, the network is optimized specifically for each wave function under consideration. In applications where the propagation of classical trajectories is the computational bottleneck, the additional cost associated with the optimization is negligible. In fact, it is even beneficial to increase the time for a proper optimization in an attempt to decrease the number of classical trajectories. Moreover, because the network is used as a variational optimizer rather than a predictive model, the usual challenges associated with neural networks, such as the construction of representative training sets and concerns about overfitting, are not relevant here. In a sense, the variational optimization adaptively identifies the set of Gaussian wave packets that best cover the relevant region of phase space explored by the input wave function.
Finally, we discuss how $N_{\text{out}}$ is determined. The optimal representation is defined by the minimum number of Gaussians used, or minimum $N_{\text{out}}$, required to achieve $F_{\text{thresh}}$, essential for reducing the number of classical trajectories. However, as end-users, we only set $K$, which is an upper limit to $N_{\text{out}}$. The algorithm starts with an initial guess of $N_{\text{out}}=N_{\text{inp}}$. If upon optimization, the threshold fidelity, $F_{\text{thresh}}$, is not achieved, then the value of $N_{\text{out}}$ is increased by a user-specified step size $\Delta N$ and the network is reoptimized. The process is repeated till either the threshold fidelity is successfully achieved or the upper limit of $K$ is reached. We refer to this procedure as an adaptive basis expansion technique, which ensures an optimal Gaussian decomposition.
@@ II.2.1 Encoding of Parameters
Now the interpretation of the output is largely similar, with the $(\vec{q},\vec{p})$ and $\gamma$, retained. The main difference comes in the $A$ matrix: the real part of $A$, $A^{\text{R}}$, must be symmetric, but the imaginary part of $A$, $A^{\text{I}}$, which encodes the wave packet width, must additionally be positive definite. Directly optimizing $A^{\text{I}}$ during training does not guarantee this property. To ensure positive definiteness by construction, we optimize a lower triangular matrix, $L$, with positive diagonal elements for each GWP. The width matrix of the $k$th Gaussian is finally reconstructed using a Cholesky decomposition, $A_{k}^{\text{I}}=L_{k}L_{k}^{T}$.
@@ II.2.2 Numerical Stabilization Strategies
VAGD, as defined so far, allows for arbitrarily broad Gaussians to be incorporated in the decomposition if required. While broad Gaussians may help minimize the number of trajectories, the local harmonic approximation is violated sooner for them than for the narrow wave packets, necessitating a more frequent decomposition to maintain accuracy. This issue can be mitigated by putting a dimension-specific, user-defined cap on the width of each Gaussian, which effectively sets a lower limit on the diagonal values of imaginary parts of $A_{k}$, such that $(A^{\text{I}}_{k})_{j,j}\geq(W_{\text{min}})_{j}$ for all dimension, $j$.
A direct implementation of a VAGD decomposition step as described till now starts from a randomly initialized network, which increases the number of optimization cycles, or epochs, required for convergence to $F_{\text{thresh}}$. Additionally, the position and the width of the wave function can fluctuate significantly during time-evolution, which makes the numerical optimization of the neural network difficult. We will see that these two difficulties are closely related.
@@
This regularization consists of two preprocessing steps. First, the wave function is recentered to remove its average position. Second, the coordinate axes are rescaled so that the wave function has variances comparable to the initial state. The neural network is trained on this preprocessed wave function. After optimization, the regularization transformation is undone to recover the physical representation.
@@ III.1 Morse Potential
We start our exploration with the dynamics of wave packets in Morse potentials. In Sec. III.1.1, we explore the convergence of VAGD-TGA to the full quantum results under different levels of anharmonicity. Subsequently in Sec. III.1.2, we study the scaling of the algorithm by looking at the dynamics under multidimensional independent Morse oscillators.
@@
where $D_{e}$ is the dissociation energy, $a$ is the width of the well, and $x_{\text{eq}}$ is the equilibrium bond distance. It is often more convenient to think in terms of the harmonic frequency at the bottom of the well, $\omega=a\sqrt{2D_{e}}$, and an anharmonicity factor, $\chi=\frac{\omega}{4D_{e}}$[1], which is zero for harmonic potentials. The parameters, $D_{e}$ and $a$ can be simply obtained by inverting the relations.
In Fig. 2, we show the overlap of the wave function evolved using the traditional TGA with the numerically exact SOFT simulations for different values of $\chi$. Following the examples in Ref. 1, the equilibrium bond length $x_{\text{eq}}=$20.95\text{\,}\mathrm{\text{ $a_{0}$ }}$\$ and $\omega=$900\text{\,}{\mathrm{cm}}^{-1}$\$. The initial wave packet is taken to be a real Gaussian centered at $q_{0}=0$, with $A=-\frac{\omega_{0}}{\hbar}$, where $\omega_{0}=$1000\text{\,}{\mathrm{cm}}^{-1}$\$. The simulations were run with a time-step of $\Delta t=$0.053\text{\,}\mathrm{fs}$\$. Consistent with our expectations, the quality of TGA deteriorates with time, and this deterioration is faster for larger values of $\chi$.
@@
Now, we explore our VAGD method. We perform the variational decomposition with $N_{\text{seg}}=50$. At each decomposition boundary, the reconstructed state is required to satisfy $F_{\text{thresh}}=0.9995$ with the propagated wavefunction. The width parameter, $W_{\text{min}}$, was set to 5.
Consider the case of $\chi=0.005$ as a first example. In Fig. 3, we examine the improvement of VAGD with increasing $K$. There is a clear and systematic improvement of the overlap with the SOFT result with the maximum number of GWPs allowed, $K$. This improvement is most clearly observed at longer times, as the quality of the semiclassical approximation gets worse. By $K\geq 6$, the VAGD-TGA becomes practically identical to SOFT. For this example, we push the limits and claim convergence at $K=8$.
Figure 4: Overlap of VAGD-TGA wave function with SOFT solution (solid, left $y$-axis) and number of VAGD trajectories (dashed, right $y$-axis) of 1D-Morse oscillator for different values of $\chi$.
On increasing anharmonicity, while $\chi=0.01$ converged with the same $N_{\text{seg}}$, $\chi=0.02$ required more frequent decompositions with $N_{\text{seg}}=25$. We show the overlap and the number of trajectories for the converged simulation in Fig. 4. The convergence gets more difficult with anharmonicity. One way to think about this is greater the anharmonicity, quicker is the breakdown of the local harmonic approximation for GWPs of the same width. Therefore, we need to perform the VAGD steps more frequently.
The number of trajectories, $N_{\text{traj}}(t)$, also shown in Fig. 4, grows with time demonstrating the build-up of the effects of anharmonicity and deviation from the basic thawed Gaussian ansatz. However, what is extremely heartening is that the number of trajectories required is relatively small in all the cases. This becomes especially important when running ab initio dynamics where trading the cost of a better decomposition for fewer number of trajectories can prove to be extremely lucrative. Moreover, depending on the physics of the problem at hand, the actual number of trajectories may be substantially less than the upper limit $K$ provided. This is another benefit of the adaptive algorithm. We emphasize that checking the overlap with the numerically exact SOFT results is by far the most stringent test of the correctness of the method. Other observables, like correlation functions, that are crucial for spectra, would converge with far fewer trajectories.